Suppose you buy 100 shares of ABC Company at $50 today and write one MAY 48 put at $3.50. If at expiry of the put, ABC’s stock price is $55, what is your total gain or loss?
Answer #1

Amount invested is = 50×100 = 5000

Premium received on put option is 3.5×100 = 350

On the date of expiry price is 55 so put option is out of money so it will not be exercised

So realised gain is 350

Unrealised gain is appreciation in stock is 5×100= 500

Total gain is 500+350= 850